# ReverseNER: A Self-Generated Example-Driven Framework for Zero-Shot Named Entity Recognition with Large Language Models

### v2

@@ 1 Introduction
The advent of Large Language Models (LLMs) has marked a significant technological leap for Natural Language Processing (NLP). Owing to extensive pre-training and parameter optimization, LLMs such as GPT-3 [2] and BERT (Bidirectional Encoder Representations from Transformers) [6] have demonstrated exceptional performance across various NLP tasks. By capturing semantics, syntax, and contextual relationships from vast corpora, these models have not only achieved breakthroughs in traditional tasks like machine translation [14], relation extraction [17], and Named Entity Recognition (NER) [16], but also exhibited strong generalization capabilities in generative tasks such as text generation [12] and dialogue generation.
Despite the impressive performance of LLMs in multiple tasks, they still face challenges in zero-shot NER, which requires models to correctly identify entity types without having seen any previously annotated data. In contrast, traditional NER methods typically rely on a large amount of annotated data for supervised learning [9]. Although LLMs possess some transfer learning capabilities, when enabling them to complete certain NER tasks without annotations, they still constantly struggle with complex tasks due to insufficient generalization abilities and imprecise understanding of domain-specific entities. These shortcomings arise from the model’s dependence on the diversity and richness of the training corpora for recognizing specific entity contexts, and when faced with new entity types or contexts, the models often struggle to match the entity categories accurately.
@@ 2.1 Named Entity Recognition
Named Entity Recognition is a fundamental task in Natural Language Processing, aimed at detecting named entities from text and categorizing them into predefined entity types. Traditional research approaches frame the NER task as a sequence labeling problem. For instance, Hammerton et al. [7] utilizes unidirectional LSTMs to obtain token-level representations, which are then fed into a softmax classifier for entity classification; Collobert et al. [3] employs CNNs to embed each input word and used Conditional Random Fields (CRFs) to decode each embedding into entity labels; Devlin et al. [6] acquires token-level representations through BERT for classification purposes; Sarzynska-Wawer et al. [13] enhances the quality of word representations by using large-scale pre-trained models; Li et al. [10] formulates the NER task as a Machine Reading Comprehension (MRC) task and further optimizes the performance of the MRC model using the dice loss function.
@@ 2.2 Applications of LLM in NER
Following this, RoBERTa, an optimized version of BERT, is introduced by Liu et al.[11], which demonstrates further improvements in NER by training on larger datasets with better hyperparameter tuning, dynamic masking, and longer training duration. RoBERTa outperformed BERT on several NER benchmarks.
Additionally, there has been an exploration of transforming NER into a generative task. Yan et al. proposed using GPT models [2] for NER by generating sentences with labeled entities instead of relying on the traditional sequence labeling format. This shift toward a generation-based approach allows for more flexible zero-shot learning settings, particularly useful in scenarios with minimal labeled data [19].
Prompt-based learning methods also gained traction for NER, especially in few-shot and zero-shot scenarios. Cui et al[4]. investigated how cloze questions and prompts could be used to leverage LLMs’ pretraining capabilities for recognizing entities with minimal task-specific training. This method shows promise in extracting entities with little to no labeled examples.
Most recently, Xie et al[18]. introduced a novel self-improving framework for zero-shot NER. Their approach utilizes an LLM to generate self-annotated datasets from an unlabeled corpus. The system then selects reliable pseudo-labeled data through self-consistency (SC) filtering strategies, and finally, this refined dataset is used for in-context learning (ICL) during inference. This self-improving strategy significantly enhances zero-shot performance on NER tasks across several benchmarks, including CoNLL03 [15] and WikiGold [1].
@@ 3.4 Entity-Level Self-Consistency Scoring
To mitigate the impact of LLM hallucinations, we perform multiple reasoning attempt rounds on the same sentence to be recognized, and introduce the self-consistency (SC) scoring mechanism. Different from the traditional Response-Level SC methods [wang2022self], we propose the Entity-Level SC Scoring method.
@@ 4.2 Overall Performance
Relative to the Vanilla zero-shot baseline, the zero-shot ReverseNER with SC demonstrates a substantial enhancement in the average F1 score (79.10 vs. 71.22), reflecting a significant improvement in the model’s capacity to generalize to unseen data without the need for additional gold labels provided in advance. Furthermore, in comparison to the few-shot methods, both with and without SC, the zero-shot ReverseNER with SC exhibits superior performance, surpassing the few-shot SC approach (77.25 average) by 1.85 points.

### v3

@@ 1 Introduction
The advent of Large Language Models (LLMs) has marked a significant technological leap for Natural Language Processing (NLP). Owing to extensive pre-training and parameter optimization, LLMs such as GPT-3 [3] and BERT (Bidirectional Encoder Representations from Transformers) [6] have demonstrated exceptional performance across various NLP tasks. By capturing semantics, syntax, and contextual relationships from vast corpora, these models have not only achieved breakthroughs in traditional tasks like machine translation [16], relation extraction [10.1145/3357384.3358119], and Named Entity Recognition (NER) [15], but also exhibited strong generalization capabilities in generative tasks such as text generation [13] and dialogue generation.
Despite the impressive performance of LLMs in multiple tasks, they still face challenges in zero-shot NER, which requires models to correctly identify entity types without having seen any previously annotated data. In contrast, traditional NER methods typically rely on a large amount of annotated data for supervised learning [12]. Although LLMs possess some transfer learning capabilities, when enabling them to complete certain NER tasks without annotations, they still constantly struggle with complex tasks due to insufficient generalization abilities and imprecise understanding of domain-specific entities. These shortcomings arise from the model’s dependence on the diversity and richness of the training corpora for recognizing specific entity contexts. When faced with new entity types or contexts, the models often struggle to match the entity categories accurately.
@@ 2.1 Named Entity Recognition
Named Entity Recognition is a fundamental task in Natural Language Processing, aimed at detecting named entities from text and categorizing them into predefined entity types. Traditional research approaches frame the NER task as a sequence labeling problem. For instance, Hammerton et al. [10] utilize unidirectional LSTMs to obtain token-level representations, which are then fed into a softmax classifier for entity classification; Collobert et al. [5] employ CNNs to embed each input word and used Conditional Random Fields (CRFs) to decode each embedding into entity labels; Sarzynska-Wawer et al.[14] enhance the quality of word representations by using large-scale pre-trained models; Li et al. [19] formulate the NER task as a Machine Reading Comprehension (MRC) task and further optimizes the performance of the MRC model using the dice loss function; Asghari et al. [1] propose BINER, a low-cost biomedical named entity recognition model that combines Bidirectional LSTM and Conditional Random Fields, achieving competitive performance with limited resources.
@@ 2.2 Applications of LLM in NER
Following this, Zhang et al.[22] introduced a finance-specific Chinese NER approach by integrating pre-trained language models with boundary detection and document-level representations, enhancing the recognition of complex entities. Fan et al.[9] improved domain-specific NER using a multi-attention DNN that combines embeddings, CNNs, and attention, achieving advanced results on biomedical datasets.
Additionally, there has been an exploration of transforming NER into a generative task. Wang et al. [15] propose a method that leverages GPT models for NER by outputting sentences with labeled entities, rather than relying on the traditional sequence labeling approach. This shift to a generation-based framework enables more flexible zero-shot learning, making it particularly advantageous in scenarios with limited labeled data. LlmNER [8], introduced by Villena et al., harnesses large language models to enable efficient zero-shot and few-shot NER.
Prompt-based learning methods also gain NER traction, especially in few-shot and zero-shot scenarios. Cheng et al. [4] enhance few-shot NER by structuring prompts into task definition, demonstration, and format, using feedback and error integration for improved performance. Ashok et al.[7] introduce PromptNER, an approach that utilizes cloze-style prompts and entity definitions to leverage the pretraining capabilities of LLMs for recognizing entities with minimal task-specific training. This method shows promise in extracting entities with little to no labeled examples.
Most recently, Xie et al. [17] introduce a self-improving framework for zero-shot NER. Their approach utilizes an LLM to generate self-annotated datasets from an unlabeled corpus. The system then selects reliable pseudo-labeled data through self-consistency (SC) filtering strategies, and finally, this refined dataset is used for in-context learning (ICL) during inference. This self-improving strategy enhances zero-shot performance on NER tasks across several benchmarks, including CoNLL03 [18] and WikiGold [2].
@@ 3.4 Entity-Level Self-Consistency Scoring
To mitigate the impact of LLM hallucinations, we perform multiple reasoning attempt rounds on the same sentence to be recognized, and introduce the self-consistency (SC) scoring mechanism. Different from the traditional Response-Level SC methods [20], we propose the Entity-Level SC Scoring method.
@@ 4.2 Overall Performance
Relative to the Vanilla zero-shot baseline, the zero-shot ReverseNER with SC demonstrates a substantial enhancement in the average F1 score (79.10 vs. 70.97), reflecting a significant improvement in the model’s capacity to generalize to unseen data without the need for additional gold labels provided in advance. Furthermore, in comparison to the few-shot methods, both with and without SC, the zero-shot ReverseNER with SC exhibits superior performance, surpassing the few-shot SC approach (77.25 average) by 1.85 points.